2. The random variable X has density ke-lal. Find P(|X| + |X – 31 < 5). A. .716 B. .235 C. .807 D. .344
Answer #1

ANSWER :  

The random variance X~ denotes Componential distance.

F_{X}(x) =\frac{1}{2}e^{-|x|}

|x|+|x-3|=\begin{cases} x+(x-3) \text{ if } x>3 \\ x+(3-x)\text{ if }0< x<3\\ -x+(3-x) \text{ if } -3<x<0\\ -x -(3-x) \text{ if } x<-3 \end{cases}

=\begin{cases} 2x-3\text{ if } x>3 \\ 3\, \, \, \, \, \, \, \, \text{ if }0< x<3\\ 3-2x\text{ if } -3<x<0\\ 3-2x \text{ if } x<-3 \end{cases}

=\begin{cases} 2x-3\text{ if } x>3 \\ 3\, \, \, \, \, \, \, \, \, \text{ if } 0<x<3 \\ 3-2x \text{ if } x<0 \end{cases}

P(|x|+|x-3|\leq 5=\begin{cases} P(2x-3\leq 5)\text{ if } x>3 \\ P(3\leq 5)\, \, \, \, \, \, \, \, \, \text{ if } 0<x<3 \\ P(3-2x\leq 5) \text{ if } x<0 \end{cases}

  =\begin{cases} P(x\leq 4)\text{ if } x>3 \\ 1\, \, \, \, \, \, \, \, \, \text{ if } 0<x<3 \\ P(x\geq -1) \text{ if } x<0 \end{cases}

\therefore P\left ( |x|+|x-3| \right )

= P ( 3<x<4)+P ( 0<x<3)+P ( -1<x<0)= \int_{3}^{4}\frac{1}{2}e^{-x}dx +\int_{3}^{0}\frac{1}{2}e^{-x}dx+\int_{-1}^{0}\frac{1}{2}e^{-(-x)}dx

= \int_{3}^{4}\frac{1}{2}e^{-x}dx +\int_{3}^{0}\frac{1}{2}e^{-x}dx+\int_{-1}^{0}\frac{1}{2}e^{x}dx

= \int_{0}^{4}\frac{1}{2}e^{-x}dx +\int_{-1}^{0}\frac{1}{2}e^{x}dx

=\left [-\frac{1}{2}e^{-x} \right ] ^{4}_{0}+\left [\frac{1}{2}e^{x} \right ] ^{0}_{-1}

=\frac{1}{2}-\frac{1}{2}e^{-4} +\frac{1}{2}-\frac{1}{2}e^{-1}

=1-\frac{1}{2}(e^{-4} +e^{-1})

= 0.80690296

\simeq 0.807\,(c)